Company ABC has sales= € 56225, costs= € 31239, other expenses= € 4500, depreciation expense= € 6200, interest expense= € 5000 and the tax rate 35%.

What is the net income?
Answer #1

The following table shows the calculations

Calculations Particulars Amount
A Sales          56,225
B Costs          31,239
C Other expenses            4,500
D Depreciation            6,200
E = A-B-C-D Earnings before Interest and Taxes          14,286
F Interest Expense            5,000
G = E - F Earnings Before Taxes            9,286
H = G x 35% Taxes            3,250
I = G - H Net Income            6,036